OpenStudy (anonymous):

Integral of (cos x/ sin x) from pi/4 to pi/2

OpenStudy (anonymous):

\[\int\limits_{\pi/4}^{\pi/2} (\cos x / \sin x)\]

OpenStudy (anonymous):

What is the antiderivative of cotanget?

OpenStudy (anonymous):

u - sub for this one. try \[u=\sin(x), du=-\cos(x)dx,-du=\cos(x)dx\] and you are just about done

OpenStudy (anonymous):

Satellite73 is on a better track, follow that advice.

OpenStudy (anonymous):

so the answer is -ln abs value of sin x

OpenStudy (anonymous):

and while you are at it, might as well make \[u(\frac{\pi}{4})=\frac{\sqrt{2}}{2}\] \[u(\frac{\pi}{2})=1\] so you don't have to change back when you integrate

OpenStudy (anonymous):

get \[-\int_{\frac{\sqrt{2}}{2}}^1\frac{du}{u}\]

OpenStudy (anonymous):

i thought that was that after u integrate du/u you get -ln sinx and then u evaluate that at both pi/4 - pi/2